# Topic-based Watermarks for LLM-Generated Text

### v1

@@ Abstract.
Large Language Models (LLMs), Watermarking, Fingerprinting
@@ 1. Introduction
The rapid expansion of artificial intelligence and natural language processing (NLP) in recent years has allowed for the development of large language models (LLMs) to become a prominent aspect in academia and industry (Inc., 2024). Models such as OpenAI’s ChatGPT (OpenAI, 2024), enables capabilities for generating human-like text by understanding user input and learning from extensive amounts of data (Ray, 2023; Zhao et al., 2023b). As more users are drawn to these systems, the amount of LLM-generated text over the internet increases, causing potential misinformation, copyright, and plagiarism issues (Rillig et al., 2023; Kalla and Kuraku, 2023). These concerns result in the need for tools that can differentiate LLM- and human-generated material.
@@ 3. Related Work
In this section, we discuss, fingerprinting methodologies (that do not target LLMs), to the best of our knowledge, the first watermarking scheme for LLMs and extensions of robust watermarking schemes.
3.1. Fingerprinting
Fingerprinting is a technique used to identify content by customizing (fingerprinting) every instance of a distributed copy of the content. By doing this, if the content is distributed without authorization, the original source of the data leak can be identified (Bailey, 2007). The method is analogous to how human fingerprints are used to identify individuals. Research concerning the robustness of fingerprinting methodologies exists in a variety of fields to allow claim to copyright and to identify sources of data breaches. Prior research by Ji et al. (Ji et al., 2022) focus on robust fingerprinting methods of genomic databases. They guarantee liability when sharing genomic data, specifically mitigating correlation attacks (an attack aiming to detect and distort the added fingerprint) against malicious database recipients by changing or fingerprinting database entries. Multiple fingerprinting schemes have shown robustness against different threat models (Ji et al., 2022; Jiang et al., 2022; Ji et al., 2023), however a notable limitation of many fingerprint approaches is their inapplicability within the content of an LLM.
3.2. Watermarking Algorithms for LLMs (Kirchenbauer et al., 2023)
@@ 3.1. Fingerprinting
The generation of the $R_{h}$ and $R_{s}$ lists utilize the input prompt and a prior sequence of known tokens containing the first $t-1$ tokens already produced by the LLM to either generate probabilities (probability of each word over the vocabulary of the LLM) or logits (discrete probabilities over the vocabulary of the LLM) depending on the type of $R$ list. By utilizing the last known token over the input sequence, a hash is computed to seed a random number generator which is then used to randomly partition the vocabulary the into $G$ and $R$ lists.
The difference between the $R_{s}$ and $R_{h}$ lists is due to the $R_{s}$ list utilizing the addition of a constant “hardness” parameter to each of the $G$ list logits. The hardness parameter allows the sampling distribution for the next token in the generated sequence to strongly favor the $G$ list for high entropy tokens. High entropy tokens are a sequence of tokens that are good choices for the next token in the generated sequence, while low entropy tokens consist of one or few good choices of tokens to choose from as the next token in the generated output. The detection of the watermark for both $R$ lists are similar and it is done by checking each token in a target text sequence and examine for violations in the $R$ list.
@@ 3.3. Robust Watermarking
Limitations considering robustness of watermarking schemes has led to an influx of research for improving the constraints of these initial algorithms (Smith et al., 2023; Liu et al., 2024, 2023; Tang et al., 2023; Christ et al., 2023; Li et al., 2024; Zhao et al., 2023a). The first watermarking algorithm extends its robustness, describing a robust private watermarking algorithm (Kirchenbauer et al., 2023). The robustness extension of the watermark involves the same generative process of previously mentioned watermarking schemes with the addition of a single secret key or multiple secret keys. The secret key(s) enhances the security and robustness on the watermark by generating unique tokens for the $R$ list utilizing a pseudorandom function. The additions of the key(s) further address the limitation of reproducibility concerns for the $R$ and $G$ lists.
@@
Continuation of work dedicated to the lack of attack robustness and security robustness, Liu et al. (Liu et al., 2024) examine a proposed watermarking algorithm to enhance attack robustness on different semantically invariant perturbations (e.g., text modifications, text alterations), embedding watermarkings based on semantics of the entire prior text instead of a fixed number of prior tokens. Analysis of their proposed semantic invariant watermarking scheme maintained robustness against attacks schemes such as text paraphrasing (rewriting the watermarked text while preserving its original meaning) and synonym substitution (certain watermarked words are replaced with their respective synonyms). Christ et al. (Christ et al., 2023) build on secret key utilization proposing an undetectable watermarking scheme utilizing cryptographic one-way functions making it computationally infeasible to distinguish a watermarked output compared to other non-watermarked text. The watermarking scheme embeds a secret in the output text to which it does not affect the text quality generated by the LLM. The detection algorithm of the undetectable watermarking scheme identifies the output of the watermarked LLM without any information other than the output text sequence and the secret key.
@@
While research in robustness of watermarking schemes for LLMs is quickly advancing, to our knowledge, there is little work concerning actual commercialized scenarios in which a usable watermarking detection scheme is feasible. Li et al. (Li et al., 2024), proposes a watermark scheme usable in the commercial settings, introducing different backdoor data paradigms which integrate the watermark to the LLM during the tuning process. The motivation of the authors is to guarantee the copyright of a customized LLM from a business perspective comparable to prior fingerprinting schemes concerning the copyright of genomic databases (Ji et al., 2022). Prior research has also assumed black-box detection approaches where the internal workings of the model are undisclosed to the user. Limitations for black-box detection schemes revolve around data biases where the data collection of LLMs can introduce biases impacting performance and the ability to generalize across different context of the detection mechanism, confidence calibration affecting the overall evaluation metrics of the LLM, indicating the reliability and trustworthiness of the LLM, and the lack of adaptability to advance the learning of the LLM (Tang et al., 2023). Our proposed watermarking methodology differentiates from existing literature with an approach to store the $G$ and $R$ lists tailored to a specified topic of the text input sequence reducing computation while still providing robustness from an attack model perspective.
@@ 4.1. Threat Model
Embedded watermarks in an LLM’s generated output are subject to various types of attacks summarized in the following. Note that all considered threat models are those where the motivation of an attacker is to distort or remove the watermark from the target text sequence to avoid detection. However, we hypothesize manipulation to the text output sequence will degrade the overall quality of text and the topic accuracy. To assess the impact of manipulations on the text quality, we utilize the evaluation metric, perplexity seen in prior research (Kirchenbauer et al., 2023) to measure how well a probability model predicts a sampled token. The perplexity metric is relevant for quantifying the degradation in text quality due to unexpected tokens as a result of efforts to evade watermark detection.
@@
While we examine a range of real-life applications for watermarking, we specifically highlight the submission of coursework within academic environments as an example. This scenario serves to illustrate the broader relevance and necessity of watermarking techniques, without confining the discussion solely to academia. In this context, the quality of manipulated text is determined by assessing its coherence, measured by the evaluation metric of perplexity. It also examines whether the text sequence accurately addresses and answers specific questions related to the coursework. This approach highlights the importance of both integrity and accuracy in an academic context. We assume that, in order to remain undetectable to the detection mechanism, a malicious user will manually change tokens while preserving text quality in relation to the specified topic(s). However, in the case of coursework submissions, preserving the quality extends beyond topic relevance to include precision and accuracy of the responses of the posed questions in the coursework submission. Modeling of this additional quality is yet to be realized, but addresses the importance of general and context-specific concerns in watermark detection.
@@
Potential attacks concern text insertion, substitution, and deletion in the generated output text sequence at the character, word, sentence, and multi-model level (Kirchenbauer et al., 2023; Goyal et al., 2023). Insertion is the addition or insertion of tokens into the text sequence. With the watermarked text encompassing significantly more $G$ listed tokens, the addition of new tokens may violate the $R$ list, increasing the count of $R$ listed tokens respective to the $G$ listed tokens. We assume the attacker will need to insert a significant amount of $R$ listed tokens to decrease detection power of the watermark to classify the text as human-generated. In the context of coursework submission, given a 1-page LLM output for an assignment, a student can add another one or more subsequent pages of text (human-generated) to decrease detection power. However, this involves a trade-off where if the student is able to write one or more pages worth of content manually to remove the watermark, there would be no need to use LLM-generated text at all. The scenario of insertion attacks is realistic but the scenario in which an attacker would need to manually generate the same or more content than the LLM-generated output is unrealistic, hence we expect a reasonable level of insertion tokens during our evaluation. Substitution of tokens, involve swapping tokens in the generated output text. For example, an attacker may substitute tokens in the text sequence with their respective synonym. The motivation of the attacker in an instance of substitution is to remove $G$ listed tokens and add $R$ listed tokens to degrade or remove the watermark. Deletion involves deleting tokens from a generated output sequence. If a significant amount of $G$ listed tokens are deleted, the amount of $R$ tokens would increase however, if only tokens are deleted from a text sequence, there exists a potential trade-off between the text quality and detection accuracy. We would also expect a reasonable level of deletion to exist in the evaluation phase of our research. Token insertion, manipulation, and deletion are used in tandem to distort or remove the watermark of an LLM-generated output. In the rest of this section, we discuss considered threat models an attacker can perform against the LLM in our evaluation.
@@ 4.1.1. Baseline Attack
Baseline attacks consist of the insertion, substitution, and deletion of text for a given output sequence. The attacker selects a single or a combination of techniques (insertion, substitution, and deletion) with the objective to diminish detection accuracy. In doing so, this approach may lead to a compromise in the text quality of the manipulated sequence.
@@ 4.1.2. Paraphrasing Attack
  • Paraphrasing with the Original LLM: The attacker generates a watermarked text sequence by feeding the LLM an input prompt. We assume the attacker does not manually alter any tokens in the output. The attacker subsequently supplies the watermarked output into the identical LLM which originally produced it, along with a paraphrase command (for example, “rephrase this text prompt”). The proposed watermarking scheme is robust against paraphrasing through the original LLM due to the nature of topics that form the basis of the watermarking process. For instance, if the attacker’s initial prompt contained the generated topic of “sports”, any paraphrased output would likely maintain “sports” as the topic, ensuring the watermark for the original and paraphrased outputs are constant.
  • Manual Paraphrasing: The attacker generates an output text sequence by the watermarked LLM from a given input prompt. The paraphrasing is conducted manually by the attacker without the aid of external tools. This manual paraphrasing approach is considered for shorter sequences of text, as longer sequences would necessitate more extensive intervention from the attacker. The topic-based watermarking scheme proposed is particularly effective against manual paraphrasing, especially for lengthier text sequences, which would require significant modification to mislead the detection mechanism into incorrectly classifying the LLM-generated text for human-generated text. The robustness of the proposed watermarking scheme is further enhanced by the potential deterioration in text quality and the possible change in the overall topics of the paraphrased text sequence.
@@ 4.1.4. Discrete Alteration Attack
An attacker’s motivation behind this attack is to induce alterations in tokens subsequently introducing misspelling and grammar errors. These manipulations can be executed through insertion or deletion of single or multiple characters, or even entire stings of characters within the specified output sequence, with the aim of diminishing the effectiveness of the watermark in detection. We consider an attacker which seeks to undermine the integrity of a watermarked text sequence without completely altering the text. By inserting or deleting characters within certain tokens, the attacker can modify the text to introduce errors, while minor, can impact the detectability of the watermark. For example, altering the word “watermarked” to “wat3rmark3d” through character substitution introducing misspelling by complicating the process of watermark recognition. The proposed topic-based watermark addresses current watermarking algorithm limitations through the implementation of the generalized topic scheme. Even in the presence of minor spelling or grammatical errors, the overarching topic of the output remains unaffected. This ensures that the watermark’s integrity is maintained, showcasing the robustness of the proposed approach against textual alterations.
@@ 4.1.5. Collusion Attack
A collusion attack begins with an original text sequence that has been processed by an LLM, which incorporates a watermarking scheme based on topic relevance. We assume that all LLMs have a topic-based watermarking scheme implemented for output text generation. The attacker then feeds this watermarked output into a different LLM, aiming to alter the content of the text through insertion, manipulation, or deletion. This process might modify the topics extracted from the text, depending on the extraction methods and the defined list pairs of general and specific topics ($G$ and $R$). Should the second LLM employ its detection mechanism for classification purposes, it would label the output as “LLM-generated”, effectively countering collusion attacks by recognizing the manipulation. However the introduction of multiple LLMs into this process disrupts the correlation between the $G$ and $R$ list pairs, rendering the detection of the text sequence challenging and potentially leading to incorrect classifications. While the use of multiple LLMs in text manipulation attacks presents a complex challenge for the detection mechanism due to the potential alteration of extracted topics and the breakdown of general and specific topic correlations, employing detection mechanisms can help identify LLM-generated content. Nonetheless, the effectiveness of these mechanisms might be compromised as the number of intervening LLMs increases, highlighting the need for advanced strategies in watermarking and detection to mitigate such threats.

### v2

@@ Abstract.
Large Language Models (LLMs), Natural Language Processing (NLP), Watermarking
@@ 1. Introduction
The rapid expansion of artificial intelligence (AI) and natural language processing (NLP) in recent years has allowed for the development of large language models (LLMs) to become a prominent aspect in academia and industry (Inc., 2024). Models such as OpenAI’s ChatGPT (OpenAI, 2024), enables capabilities for generating human-like text by understanding user input and learning from extensive amounts of data (Ray, 2023; Zhao et al., 2023b). As more users are drawn to these systems, the amount of LLM-generated text over the internet increases, causing potential misinformation, copyright, and plagiarism issues (Rillig et al., 2023; Kalla and Kuraku, 2023). These concerns result in the need for tools that can differentiate LLM- and human-generated material.
@@ 2. Background
Watermarking is a technique which involves embedding hidden markers or patterns within data to assert copyright, verify authenticity, and identity sources of data breaches. Research concerning the robustness of watermarking methodologies exists in a variety of fields. In prior work, Ji et al. (Ji et al., 2022) have focused on robust watermarking methods for genomic databases, ensuring accountability when sharing genomic data. Their research specifically addresses the mitigation of correlation attacks (an attack aiming to detect and distort the added watermark) against malicious database recipients by changing or watermarking database entries. While multiple watermarking schemes have demonstrated robustness against various threat models (Ji et al., 2022; Jiang et al., 2022; Ji et al., 2023), theses approaches have limited applicability within the context of LLMs. This limitation has facilitated the adoption of watermarking concepts to the field of generative AI, including areas such as NLP and LLMs. In the context of LLMs, watermarkings can be applied either to the data processed by the LLM or to the outputs they generate. These watermarking schemes are created to be invisible to human individuals, but identified through algorithmic detection mechanisms, enhancing the security and verification of generated text sequences.
@@ 3. Related Work
In this section, we explore the initial watermarking schemes developed for LLMs and examine enhancements aimed at increasing the robustness of these watermarking strategies.
3.1. Watermarking Algorithms for LLMs
@@
The generation of the $R_{h}$ and $R_{s}$ lists involve utilizing the initial input prompt and a sequence the first tokens $t-1$ that the LLM has already produced. The two $R$ lists are designed to either generate probabilities (which represent the likelihood for each token within the LLM’s vocabulary) or logits (which are discrete probabilities over the LLM’s vocabulary), depending on the specific $R$ list ($R_{h}$ or $R_{s}$) being used. By utilizing the last known token over the input sequence, a hash is computed to seed a random number generator which is then used to randomly partition the vocabulary the into $G$ and $R$ lists.
The difference between the $R_{s}$ and $R_{h}$ lists is because the $R_{s}$ list utilizes the addition of a constant “hardness” parameter to each of the $G$ list logits. The hardness parameter allows the sampling distribution for the next token in the generated sequence to strongly favor the $G$ list for high entropy tokens. High entropy tokens are a sequence of tokens that are good choices for the next token in the generated sequence, while low entropy tokens consist of one or few good choices of tokens to choose from as the next token in the generated output. The detection of the watermark for both $R$ lists are similar and it is done by checking each token in a target text sequence and examine for violations in the $R$ list.
@@ 3.2. Robust Watermarking
Limitations considering robustness of watermarking schemes has led to an influx of research for robust watermarking algorithms (Smith et al., 2023; Liu et al., 2024, 2023; Tang et al., 2023; Christ et al., 2023; Li et al., 2024; Zhao et al., 2023a). The first watermarking algorithm by Kirchenbauer et al. (Kirchenbauer et al., 2023), to extend its robustness, describes a robust private watermarking algorithm. The robustness extension involves employing the same generative process of their previously mentioned watermarking scheme, with the addition of a single secret key or multiple secret keys. The secret key(s) enhances the security and robustness on the watermark by generating unique tokens for the $R$ list utilizing a pseudorandom function. The additions of the key(s) further address the limitation of reproducibility concerns for the $R$ and $G$ lists.
@@
Motivated due to the lack of robustness in existing watermarking algorithms, Liu et al. (Liu et al., 2024) propose a watermarking algorithm to enhance attack robustness on different semantically invariant perturbations (e.g., text modifications, text alterations). They achieve this by embedding watermark based on semantics of the entire prior text instead of a fixed number of prior tokens. Their scheme maintain robustness against attacks such as text paraphrasing (rewriting the watermarked text while preserving its original meaning) and synonym substitution (certain watermarked words are replaced with their respective synonyms). Christ et al. (Christ et al., 2023) propose an undetectable watermarking scheme utilizing cryptographic one-way functions, making it computationally infeasible to distinguish a watermarked output compared to other non-watermarked text. The watermarking scheme embeds a secret in the output text to which it does not affect the text quality generated by the LLM. The detection algorithm of the undetectable watermarking scheme identifies the output of the watermarked LLM without any information other than the output text sequence and the secret key.
@@
While research in robustness of watermarking schemes for LLMs is quickly advancing, to our knowledge, there is little work concerning actual commercialized scenarios in which a usable watermarking detection scheme is feasible. Li et al. (Li et al., 2024), propose a watermarking scheme that may be practical in the commercial settings, introducing different backdoor data paradigms which integrate the watermark to the LLM during the tuning process. The motivation of the authors is to guarantee the copyright of a customized LLM from a business perspective comparable to prior watermarking schemes concerning the copyright of genomic databases (Ji et al., 2022). Li et al. focus on the copyright of an LLM in an internal company environment instead of differentiating between human- and LLM-generated text sequences. Prior research has also assumed black-box detection approaches where the internal workings of the model are undisclosed to the user. Limitations for black-box detection schemes revolve around data biases where the data collection of LLMs can introduce biases impacting performance and the ability to generalize across different context of the detection mechanism, confidence calibration affecting the overall evaluation metrics of the LLM, indicating the reliability and trustworthiness of the LLM, and the lack of adaptability to advance the learning of the LLM (Tang et al., 2023). Our proposed watermarking methodology differentiates from existing literature by enhancing practicality and robustness against attacks, specifically through our approach to store $G$ and $R$ list pairs tailored to topics of the input text sequence. Aligning with current studies, we utilize list pairs to elicit specificity and robustness in watermarking LLM-generated text, ensuring the watermark remains contextually relevant and minimally intrusive. This adaptation not only preserves text quality but also complicates watermark removal, increasing security against attacks and reducing false positives in watermark detection. Our proposed watermarking scheme further seeks to reduce computational demands and improve resistance toward known threat models, offering a practical solution for watermarking LLM-generated text.
@@ 4.1. Threat Model
Embedded watermarks in an LLM’s generated output are subject to various types of attacks summarized in the following. Note that all considered threat models are those where the motivation of an attacker is to distort or remove the watermark from the target text sequence to avoid detection. On the other hand, distorting the text output sequence will degrade the overall quality of text, and hence the topic of the corresponding text. To assess the impact of manipulations on the text quality, we propose to utilize the evaluation metric, perplexity, seen in prior research (Kirchenbauer et al., 2023) to measure how well a probability model predicts a sampled token. The perplexity metric is relevant for quantifying the degradation in text quality due to unexpected tokens as a result of efforts to evade watermark detection.
@@
While we examine a range of real-life applications for watermarking, we specifically stick to the generation and submission of coursework in academia as an illustrative example. This scenario serves to illustrate the broader relevance and necessity of watermarking techniques, without confining the discussion solely to academia. In this context, the quality of manipulated text is determined by assessing its coherence, measured by the evaluation metric of perplexity. This evaluation approach also examines whether the text sequence accurately addresses and answers specific questions related to the coursework. This approach highlights the importance of both integrity and accuracy in an academic context. We assume that, in order to remain undetectable to the detection mechanism, a malicious user will manually change tokens while aiming to preserve text quality in relation to the specified topic(s). However, in the case of coursework submissions, preserving the quality extends beyond topic relevance to include precision and accuracy of the responses of the posed questions in the coursework submission. Modeling of this additional quality is yet to be realized, but addresses the importance of general and context-specific concerns in watermark detection.
@@
Potential tools of an attacker include text insertion, substitution, and deletion in the generated output text sequence at the character, word, sentence, and multi-model (combination of characters, words, or sentences) level (Kirchenbauer et al., 2023; Goyal et al., 2023). Insertion is the addition or insertion of tokens into the text sequence. With the watermarked text encompassing significantly more $G$ listed tokens, the addition of new tokens may violate the $R$ list, increasing the count of $R$ listed tokens with respect to the $G$ listed tokens. However, the attacker will need to insert a significant amount of $R$ listed tokens to decrease detection power of the watermark to classify the text as human-generated. For instance, in the context of coursework submission, given a 1-page LLM output for an assignment, a student can add additional (human-generated) pages to decrease the detection power. However, this involves a trade-off: if the student is able to write one or more pages worth of content manually to remove the watermark, there would be no need to use LLM-generated text at all. The scenario of insertion attacks is realistic, but the fact that an attacker would need to manually generate the same or more content than the LLM-generated output is unrealistic, hence we expect a reasonable amount of insertion (i.e., manually generated tokens) as a part of the insertion attack. Substitution of tokens, involve swapping tokens in the generated output text. For example, an attacker may substitute tokens in the text sequence with their respective synonym. The motivation of the attacker in an instance of substitution is to remove $G$ listed tokens and add $R$ listed tokens to degrade or remove the watermark. Deletion involves deleting tokens from a generated output sequence. If a significant amount of $G$ listed tokens are deleted, the amount of $R$ tokens would increase however, if only tokens are deleted from a text sequence, there exists a potential trade-off between the text quality and detection accuracy. Similar to insertion, we also expect a reasonable level of deletion in a rational attack scenario. Token insertion, manipulation, and deletion are used in tandem to distort or remove the watermark of an LLM-generated output. An attacker can use a combination of these aforementioned tools to generate its attack against the watermarking scheme. In the rest of this section, we discuss considered threat models an attacker can perform against the LLM in our evaluation.
@@ 4.1.1. Baseline Attack
Baseline attack consists of the insertion, substitution, and deletion of text for a given output sequence. The attacker selects a single or a combination of techniques (insertion, substitution, and deletion) with the objective to diminish detection accuracy. In doing so, this approach may lead to a compromise in the text quality of the manipulated sequence.
@@ 4.1.2. Paraphrasing Attack
  • Paraphrasing using the original LLM: The attacker generates a watermarked text sequence by feeding the LLM an input prompt. The attacker does not manually alter any tokens in the output. Instead, the attacker subsequently supplies the watermarked output into the identical LLM which originally produced it, along with a paraphrase command (i.e., “rephrase this text prompt”). The proposed watermarking scheme is robust against paraphrasing using the original LLM due to the nature of topics that form the basis of the watermarking process. For instance, if the attacker’s initial prompt leads to an output about “sports” (in terms of its topic), any paraphrased output would maintain the same topic, ensuring the watermark (on the paraphrased output) can still be generated and detected in the same way.
  • Manual Paraphrasing: This time, the paraphrasing is conducted manually by the attacker without the aid of external tools. This manual paraphrasing approach is considered for shorter sequences of text, as longer sequences would necessitate more extensive intervention from the attacker. The proposed topic-based watermarking scheme is particularly effective against manual paraphrasing, especially for lengthier text sequences, which would require significant modification to mislead the detection mechanism into incorrectly classifying the LLM-generated text for human-generated text. The robustness of the proposed watermarking scheme is further enhanced by the potential deterioration in text quality and the possible change in the overall topics of the paraphrased text sequence.
@@ 4.1.4. Discrete Alteration Attack
An attacker’s motivation behind this attack is to induce alterations in tokens, subsequently introducing misspelling and grammar errors. These manipulations can be executed through insertion or deletion of single or multiple characters, or even entire stings of characters within the specified output sequence, with the aim of diminishing the effectiveness of the watermark in detection. We consider an attacker which seeks to undermine the integrity of a watermarked text sequence without completely altering the text. By inserting or deleting characters within certain tokens, the attacker can modify the text to introduce errors, while minor, can impact the detectability of the watermark. For example, altering the word “watermarked” to “wat3rmark3d” through character substitution introducing misspelling by complicating the process of watermark recognition. The proposed topic-based watermarking addresses current watermarking algorithm limitations through the implementation of the generalized topic scheme. Even in the presence of minor spelling or grammatical errors, the overarching topic of the output remains unaffected. This ensures that the watermark’s integrity is maintained, showcasing the robustness of the proposed approach against textual alterations.
@@ 4.1.5. Collusion Attack
A collusion attack begins with an original text sequence that has been produced by an LLM, which incorporates a watermarking scheme based on topic relevance. Ideally, we assume that all existing LLMs have a topic-based watermarking scheme implemented for output text generation. [^1] Without this assumption, all watermarking schemes are vulnerable with the help of an LLM that does not add a watermark to its output. The attacker then solicits outputs to the same prompt from several LLMs, aiming to blend these multiple responses int a single output through insertion, manipulation, or deletion. This process might modify the topics extracted from the text, depending on the extraction methods and the defined list pairs of general and specific topics ($G$ and $R$). Should any of the LLMs employ its detection mechanism for classification purposes, it would label the mixed outputs as “LLM-generated”, effectively countering collusion attacks by recognizing the manipulation. However, the involvement of multiple LLMs in this process disrupts the correlation between the $G$ and $R$ list pairs, rendering the detection of the text sequence challenging and potentially leading to incorrect classifications. While the use of multiple LLMs in text manipulation attacks presents a complex challenge for the detection mechanism due to the potential alteration of extracted topics and the breakdown of general and specific topic correlations, employing detection mechanisms can help identify LLM-generated content. Nonetheless, the effectiveness of these mechanisms might be compromised as the number of intervening LLMs increases, highlighting the need for advanced strategies in watermarking and detection to mitigate such threats.